### v1

# Real-time Scattering in $\mathbf{\phi^{4}}$ Theory using Matrix Product States

Abstract
We investigate the critical behavior and real-time scattering dynamics of the interacting $\phi^{4}$ quantum field theory in $(1+1)$ dimensions using uniform matrix product states and the time-dependent variational principle. A finite-entanglement scaling analysis at $\lambda=0.8$ bounds the critical mass-squared to $\mu_{c}^{2}\in[-0.3190,-0.3185]$ and provides a quantitative map of the symmetric, near-critical, weakly broken, and deeply broken regimes. Using these ground states as asymptotic vacua, we simulate two-particle collisions in a sandwich geometry and extract the elastic scattering probability $P_{11\to 11}(E)$ and Wigner time delay $\Delta t(E)$ following the prescription of Jha et al. [1]. We find strongly inelastic scattering in the symmetric phase ($P_{11\to 11}\simeq 0.63$, $\Delta t\simeq-180$ for $\mu^{2}=0.2$), almost perfectly elastic collisions in the spontaneously broken phase ($P_{11\to 11}\simeq 0.998$, $\Delta t\simeq-270$ for $\mu^{2}=-0.2$ and $P_{11\to 11}\simeq 1$, $\Delta t\simeq-177.781$ for $\mu^{2}=-0.5$), and a breakdown of the sandwich evolution precisely at the critical coupling, which provides a dynamical signature of the quantum critical point. These results demonstrate that TDVP-based uniform matrix product states can probe nonperturbative scattering and critical dynamics in lattice $\phi^{4}$ theory with controlled entanglement truncation.
I Introduction
Quantum field theory (QFT) provides the standard framework for describing relativistic many-body systems and phase transitions [2, 3, 4, 5]. A paradigmatic example is the quartic interaction model, or $\phi^{4}$ theory [6], whose action in $(1+1)$ dimensions reads
$S_{\phi^{4}}=\int d^{2}x\,\bigg[\frac{1}{2}(\partial_{\nu}\phi)(\partial^{\nu}\phi)-\bigg(\frac{1}{2}\mu^{2}\phi^{2}+\frac{\lambda}{4!}\phi^{4}\bigg)\bigg],$ (1)
where $\phi(x)$ is a real scalar field, $\mu^{2}$ is the bare mass-squared parameter, and $\lambda>0$ controls the strength of the quartic self-interaction. The $\phi^{4}$ model underlies the field-theoretic description of classical and quantum critical phenomena [7, 8, 9, 10] and provides a canonical realization of Ising universality in $(1+1)$ dimensions [11, 6].
Despite the success of perturbative techniques such as Feynman diagrams [12, 13] and renormalization-group methods [14], they offer limited access to real-time dynamics and strongly coupled regimes, leaving nonperturbative information about scattering processes typically extracted from collider experiments (e.g., the SPS, Tevatron, and LHC) rather than from first-principles calculations.
In two dimensions, complementary nonperturbative insights into scattering and resonance structure can be obtained from the analytic $S$-matrix program and integrable deformations of Ising field theory [15, 16, 17, 18, 19, 20, 21], but these approaches do not directly address real-time wave-packet dynamics in a given microscopic Hamiltonian.
In parallel, there is a growing effort to compute real-time scattering amplitudes directly on quantum computers. Algorithmic proposals for simulating scattering in scalar quantum field theories on fault-tolerant devices have been put forward in Ref. [22], and recent proof-of-principle digital quantum simulations have begun to implement scattering dynamics for scalar and Ising-like field theories using up to $\mathcal{O}(10^{2})$ qubits, W-state encodings, and Hamiltonian truncation on noisy intermediate-scale hardware [23, 24, 25].
In $(1+1)$ dimensions, tensor-network methods — in particular matrix product states (MPS) [26, 27, 28] combined with the time-dependent variational principle (TDVP) and tangent-space methods [29, 30, 31, 32] — provide a complementary route to real-time evolution in the thermodynamic limit. This approach has enabled simulations of false-vacuum collisions, Schwinger pair production, and gauge dynamics [33, 34, 35], and was recently used by Jha et al. to study real-time scattering in Ising field theory [1].
Here we extend TDVP-based uniform MPS (uMPS) methods to the interacting $\phi^{4}$ field theory with the goal of probing both its critical behavior and real-time two-particle scattering across the phase diagram. We focus on a fixed quartic coupling $\lambda=0.8$ and use finite-entanglement scaling (FES) [26] to locate the quantum critical point and characterize the adjacent phases. The FES analysis yields an estimate for the critical mass-squared $\mu_{c}^{2}$ and provides a quantitative map of the symmetric, near-critical, and spontaneously broken regimes that we subsequently explore via scattering simulations.
We discretize the continuum theory with lattice spacing $a$ and work in units $a=\hbar=c=1$. The Hamiltonian is written as $H=\sum_{i}h_{i,i+1}$ with
$h_{\mathrm{i,i+1}}=\frac{1}{2}\left(\pi_{i}^{2}+\pi_{i+1}^{2}\right)+\frac{1}{2}\left(\phi_{i}-\phi_{i+1}\right)^{2}+\frac{1}{2}\mu^{2}\left(\phi_{i}^{2}+\phi_{i+1}^{2}\right)+\frac{\lambda}{4!}\left(\phi_{i}^{4}+\phi_{i+1}^{4}\right),$ (2)
where on-site terms are split evenly across bonds so that $\sum_{i}h_{i,i+1}$ reproduces the standard discretization of $\frac{1}{2}\pi^{2}+\frac{1}{2}(\nabla\phi)^{2}+\frac{1}{2}\mu^{2}\phi^{2}+\frac{\lambda}{4!}\phi^{4}$. We take $\phi_{i},\pi_{j}]=i\delta_{ij}$, and treat $(\mu,\lambda)$ as bare lattice couplings; when needed we normal-order with respect to a reference mass to control the additive renormalization in (1+1)D.
The many-body state $|\psi\rangle$ is written as a translation-invariant uMPS
$|\psi(A)\rangle=\sum_{\{s_{i}\}}v_{L}^{\dagger}\Bigg(\prod_{n\in\mathbb{Z}}A^{s_{n}}\Bigg)v_{R}\,|\{s_{i}\}\rangle,$ (3)
where $A^{s}$ is a $D\times D$ matrix, identical on every site, $D$ is the bond dimension, and $s$ labels a $d$-dimensional truncated local Hilbert space.
Working directly in the thermodynamic limit, a finite bond dimension $D$ imposes an effective correlation length $\xi_{D}$ even at criticality. Near a conformal fixed point, the half-chain von Neumann entropy $S$ obeys the finite-entanglement scaling relation
$S\simeq\frac{c}{6}\,\log\xi_{D}+\text{const.},$ (4)
with $c$ the central charge and $\xi_{D}\propto D^{\kappa}$. This relation enables us to estimate the critical coupling $\mu_{c}^{2}$ and verify the Ising universality class directly from uMPS ground states, without changing system size. The resulting bracket for $\mu_{c}^{2}$ then serves as input for our real-time scattering simulations in the different phases.
In a second step, we use these uMPS ground states as asymptotic vacua for real-time scattering simulations in a nonuniform “sandwich” geometry. Following the prescription of Jha et al. [1], we imprint two localized wave packets with opposite momenta on top of the uniform background and evolve the resulting state using TDVP for a finite window embedded in the bulk. From the asymptotic outgoing state we extract the elastic scattering probability $P_{11\to 11}(E)$ and Wigner time delay $\Delta t(E)$, and track their behavior as the theory is tuned from the symmetric phase through the critical region into the spontaneously broken regime. This combination of finite-entanglement scaling and real-time scattering provides a unified, nonperturbative view of criticality and quasiparticle dynamics in lattice $\phi^{4}$ theory.
II Methods
Finite-entanglement scaling.— As discussed in the Introduction, a uniform MPS (uMPS) with finite bond dimension $D$ induces an effective correlation length $\xi_{D}$ and the entropy–length relation Eq.(4) near a conformal fixed point. Here we describe how we use this relation in practice to locate the critical coupling and test universality.
We parametrize the theory by $r\equiv\mu^{2}$, and denote the critical value by $r_{c}\equiv\mu_{c}^{2}$.
For a fixed quartic coupling $\lambda=0.8$ we compute translationally invariant ground states $|\psi_{D}(r)\rangle$ for a set of bond dimensions $\{D\}$ and a sweep of mass parameters $r$ across the putative critical region. For each pair $(r,D)$ we extract the correlation length $\xi_{D}(r)$ from the subleading eigenvalue of the uMPS transfer matrix and obtain the bipartite entropy $S_{D}(r)$ from the Schmidt spectrum. We then fit, at fixed $r$, $S_{D}(r)$ as a function of $\log\xi_{D}(r)$ to a linear form $S_{D}(r)\simeq a(r)\,\log\xi_{D}(r)+b(r),$ and interpret the slope as an effective central charge, $c_{\mathrm{eff}}(r)=6\,a(r)$.
The resulting $c_{\mathrm{eff}}(r)$ profile exhibits a pronounced peak as $r$ approaches criticality. We identify a narrow interval in $r$ that contains this peak and take it as a bracket for the critical coupling,
$r_{c}\in[r_{\min},r_{\max}],$ (5)
with the peak position furnishing a central estimate and variations of the fit window and $D$-range providing an uncertainty. As a minimal consistency check we also test data collapse of a local observable (e.g. the order parameter) when plotted as a function of $(r-r_{c})\,\xi_{D}^{1/\nu}$, with $\nu=1$ (Ising). This confirms that the extracted $r_{c}$ is compatible with Ising universality within our accessible bond dimensions. The FES analysis thus yields a quantitatively controlled estimate of the critical mass-squared and a map of the surrounding parameter space that we use to select representative points in the symmetric, near-critical, weakly broken, and deeply broken regimes.
Real-time scattering protocol.— To probe real-time scattering we follow the sandwich geometry introduced in Ref. [1]. For each choice of mass parameter $r$ at fixed quartic coupling $\lambda=0.8$ we take the corresponding uMPS ground state $|\Omega(r)\rangle$ as an approximation to the asymptotic vacuum. We then imprint two opposite-momentum wave packets on top of $|\Omega(r)\rangle$ at well-separated positions,
$|\Psi(0)\rangle=\hat{B}^{\dagger}_{-k}(x_{L})\,\hat{B}^{\dagger}_{+k}(x_{R})\,|\Omega(r)\rangle,$ (6)
where $\hat{B}^{\dagger}_{\pm k}(x)$ creates a localized quasiparticle with central momentum $\pm k$ and a smooth spatial profile such that the initial packets are approximately non-overlapping. The subsequent evolution $|\Psi(t)\rangle$ is obtained by integrating the TDVP equations for a non-uniform MPS defined on a finite window that contains both packets and is embedded in the uniform background $|\Omega(r)\rangle$.
We monitor the local field expectation value $\langle\phi_{n}(t)\rangle$ and its connected correlators throughout the collision. Once the outgoing packets are again well separated, we analyze the asymptotic state to extract the elastic scattering probability $P_{11\to 11}(E)$ at the corresponding center-of-mass energy $E$. In particular, we employ the diagonal projection scheme of Ref. [1], adapted to the $\phi^{4}$ Hamiltonian, which yields a numerically stable estimate of the probability that both outgoing packets contain exactly one quasiparticle. The Wigner time delay $\Delta t(E)$ is obtained from the shift in the peak position of the outgoing wave packets relative to a non-interacting reference evolution at the same energy. This protocol is applied at representative points in the symmetric, near-critical, weakly broken, and deeply broken regimes selected by the FES analysis above.
All ground-state optimizations and real-time evolutions are implemented using a modernized version of the open-source evoMPS code developed by Milsted [36], which we have extended to support the $\phi^{4}$ Hamiltonian and the scattering diagnostics used here.
III Results
Finite-entanglement scaling.— Using uMPS with bond dimensions $D=\{48,64,96\}$, we performed a coarse-grained scan of the mass-squared parameter $r\equiv\mu^{2}$ over a broad interval and identified an approximate critical region around $r\simeq-0.32$. We then carried out a refined scan in the window $r\in[-0.325,-0.310]$ using the same FES procedure. For each pair $(r,D)$ we extracted the correlation length $\xi_{D}(r)$ from the subleading eigenvalue of the uMPS transfer matrix and the bipartite entropy $S_{D}(r)$ from the Schmidt spectrum, and fitted $S_{D}(r)$ as a function of $\log\xi_{D}(r)$ to obtain the slope $a(r)$ and effective central charge $c_{\mathrm{eff}}(r)=6\,a(r)$, as described in the Methods.
(a) $c_{\text{eff}}$ vs $\mu^{2}$.
(b) $\langle\phi\rangle$ across the phase transition.
(c) $\mathcal{S}$ vs $\log\xi$.
(d) $\xi$ vs $D$
Figure 1: Finite-entanglement scaling near the critical point of the $\phi^{4}$ theory at $\lambda=0.8$. Ground-state uMPS data with finite bond dimension $D$ are used to extract the effective central charge $c_{\mathrm{eff}}(r)$, a local order parameter, and the entropy–correlation-length relation near $r_{c}=\mu_{c}^{2}$. The peak in $c_{\mathrm{eff}}(r)$ and the linear dependence $\mathcal{S}\simeq\frac{c}{6}\log\xi_{D}+\mathrm{const.}$ close to $r_{c}$ yield the bracket $\mu_{c}^{2}\in[-0.3190,-0.3185]$ and confirm Ising universality.
Across the sweep, $c_{\mathrm{eff}}(r)$ grows rapidly as $r$ approaches $r\simeq-0.319$ and exhibits a clear maximum within our grid:
$c_{\mathrm{eff}}(-0.3190)$ $=0.118\pm 0.022,$
$c_{\mathrm{eff}}(-0.3185)$ $=0.191\pm 0.102,$
$c_{\mathrm{eff}}(-0.3180)$ $=0.063\pm 0.015.$ (7)
We therefore bracket the critical point as
$r_{c}\in[-0.3190,-0.3185],$ (8)
with a central estimate near $r_{c}\approx-0.31875$. Given the modest entanglement range $D\in[48,96]$ and the relatively coarse sampling in $r$, the peak height underestimates the Ising value $c=1/2$, as expected from finite-$D$ truncation, but the trend of $c_{\mathrm{eff}}(r)$ is fully consistent with Ising universality.
For the subsequent scattering study we use this FES map to select representative couplings. In particular, we choose a symmetric-phase point $r=+0.2$, a near-critical point close to the bracketed $r_{c}$, and a weakly broken point r=0.2r=-0.2 within the spontaneously broken phase. These choices ensure that the scattering simulations sample distinct regions of the phase diagram while remaining quantitatively anchored to the FES estimate of $r_{c}$. A more deeply broken point at r=0.5r=-0.5 is also included to illustrate the strongly gapped regime.
Scattering in the $\phi^{4}$ theory.— Before studying real-time dynamics, we characterize the excitation spectrum through the single-particle and multi-particle bands shown in Fig. 2. In the symmetric phase at $\mu^{2}=+0.2$ [Fig. 2(a)], the system is gapped: the lowest excitation at $p=0$ has energy $dE\approx 1.2$, and there is a clear gap between the single-particle band and the multi-particle continuum, which onsets around $dE\approx 2.4$. This is characteristic of a stable symmetric phase, with a vacuum at $\phi=0$ and standard massive bosonic excitations.
(a) $\mu^{2}=+0.2$
(b) $\mu^{2}=-0.31875$
(c) $\mu^{2}=-0.2$
(d) $\mu^{2}=-0.5$
Figure 2: Excitation spectra of the $\phi^{4}$ theory at $\lambda=0.8$ across the phase diagram. Each panel shows the single- and multi-particle excitation energies $dE(p)$ obtained from the uMPS transfer matrix as a function of momentum $p$ for representative values of $\mu^{2}$. The gap closes near (b) and reopens with a massive single-particle band in (c),(d), providing a spectral signature of the phase transition and mass generation.
Very close to the critical point at $\mu^{2}_{c}\approx-0.31875$ [Fig. 2(b)] the spectrum becomes gapless. The mass of the fundamental particle vanishes ($m=0$), and the lowest band starts at $dE\approx 0$, while the multi-particle thresholds (at $2m$, $3m$, etc.) collapse onto one another. The spectrum is consistent with an underlying conformal field theory: the dispersion near $p=0$ is approximately linear, $E\propto|p|$, and the stacked structure of the low-lying bands suggests a tower of stable or long-lived bound states at this coupling. In the spontaneously broken phase [Figs. 2(c), 2(d)], the system is gapped again and a single-particle band re-emerges with a nonzero mass $m_{s}$. As the system moves away from criticality into the deeply broken regime (e.g., $\mu^{2}=-0.5$), this mass increases, providing a spectral signature of mass generation across the phase diagram.
To probe the nonperturbative dynamics we simulate real-time two-particle collisions across these regimes, as shown in Fig. 3. In the gapped phases, panels [3(a)], [3(c)], and [3(d)] display the characteristic diagonal “X" pattern in $\langle\phi_{n}(t)\rangle$, indicative of two localized wave packets approaching, interacting, and separating again as outgoing quasiparticles. The overall structure confirms the stable propagation of massive excitations in both the symmetric and spontaneously broken phases.
(a) $\mu^{2}=+0.2$
(b) $\mu^{2}=-0.31875$
(c) $\mu^{2}=-0.2$
(d) $\mu^{2}=-0.5$
Figure 3: Real-time two-particle scattering in the $\phi^{4}$ theory at $\lambda=0.8$, shown as space–time plots of $\langle\phi_{n}(t)\rangle$ in the sandwich geometry. Panels (a)–(d) correspond to the same values of $\mu^{2}$ as in Fig. 2. The gapped phases, (a),(c),(d), display the characteristic “X” pattern of wave packets approaching, colliding, and separating as outgoing quasiparticles, with strong inelasticity in the symmetric case (a) and almost perfectly elastic scattering in the broken phases (c),(d). At the critical point (b) the “X” pattern is absent and a slow drift of the profile indicates a breakdown of the sandwich protocol due to the diverging correlation length.
The detailed scattering outcome, however, depends strongly on the phase. In the symmetric phase [Fig. 3(a), $\mu^{2}=+0.2$], the collision is highly inelastic: the extracted elastic probability and Wigner time delay at our reference energy $E_{0}$ are
$P_{11\rightarrow 11}(E_{0})$ $\simeq 0.636,$ (9)
$\Delta t_{\mathrm{sym}}(E_{0})$ $\simeq-180.718.$ (10)
signaling a substantial cross section for particle production and other inelastic channels, as expected for a nonintegrable theory with a fundamental boson. By contrast, in the weakly broken phase [Fig. 3(c), $\mu^{2}=-0.2$] the scattering is almost perfectly elastic,
$P_{11\rightarrow 11}(E_{0})$ $\simeq 0.998,$ (11)
$\Delta t_{\mathrm{br}}(E_{0})$ $\simeq-269.720,$ (12)
with the incoming wave packets essentially re-emerging as outgoing packets with the same particle content and minimal radiation. The deeply broken point [Fig. 3(d), $\mu^{2}=-0.5$] exhibits similar qualitatively elastic behavior, $P_{11\rightarrow 11}\simeq 1$ and $\tau\simeq-177.781$, with even more clearly separated outgoing wave fronts due to the larger gap.
At the critical coupling $\mu_{c}^{2}=-0.31875$ [Fig. 3(b)], the sandwich evolution ceases to produce a well-defined scattering event. Instead of the characteristic “X”-shaped collision pattern, $\langle\phi_{n}(t)\rangle$ develops a long-wavelength drift across the entire window, and no clean separation into incoming and outgoing wave packets is observed. This is not a numerical artifact or an instability of TDVP/uMPS, but a direct consequence of criticality: the sandwich construction assumes a finite correlation length $\ell$ that separates a central scattering region from asymptotic bulk domains. At the critical point the exact ground state is gapless ($m=0$) and scale invariant, with a diverging correlation length ($\ell\to\infty$), so any finite-window implementation of this protocol inevitably violates that assumption and fails to reach asymptotic scattering states. In this sense, the breakdown of the sandwich geometry at $\mu^{2}=\mu_{c}^{2}$ reflects the underlying physics of an infinitely correlated critical state rather than a flaw of the method, and it provides a dynamical signature of the quantum critical point that is complementary to the static FES diagnostics discussed above. Closer and closer to $\mu_{c}^{2}$ the required window sizes and evolution times grow with $\ell$, so in practice the protocol becomes rapidly more expensive and ultimately impractical as the critical point is approached.
IV Summary and Discussion
We have extended TDVP-based uniform matrix product state techniques, previously applied to Ising field theory [1], to the interacting $\phi^{4}$ quantum field theory in $(1+1)$ dimensions. A finite-entanglement scaling analysis at $\lambda=0.8$ yields a controlled estimate of the critical mass-squared, $\mu_{c}^{2}\in[-0.3190,-0.3185]$, consistent with Ising universality and with the expected emergence of a gapless, conformal regime. This provides a quantitative map of the symmetric, near-critical, weakly broken, and deeply broken phases, anchoring the subsequent real-time simulations to a data-driven bracket of the critical point rather than to ad hoc parameter choices.
Using these uMPS ground states as asymptotic vacua, we simulated two-particle collisions in a sandwich geometry across the phase diagram and extracted both the elastic scattering probability $P_{11\to 11}(E)$ and the Wigner time delay $\Delta t(E)$. In the symmetric phase the collisions are strongly inelastic, with $P_{11\to 11}\approx 0.63$, reflecting the sizable phase space for particle production in a nonintegrable theory with a fundamental boson. In the spontaneously broken phase the same protocol yields almost perfectly elastic collisions, $P_{11\to 11}\approx 0.998$, indicating that the excitations behave as stable, weakly radiating quasiparticles. As the system is tuned deeper into the broken regime, the increasing gap sharpens this behavior and further suppresses inelastic channels.
At the critical coupling $\mu^{2}=\mu_{c}^{2}$ the sandwich evolution fails to produce a clean scattering event: instead of a well-defined “X” pattern in $\langle\phi_{n}(t)\rangle$, we observe a long-wavelength drift across the entire window and no clear separation into incoming and outgoing packets. This breakdown is a direct manifestation of criticality: the scattering construction relies on a finite correlation length to isolate a central interaction region from asymptotic bulk domains, an assumption that is invalidated when the mass gap closes and $\xi\to\infty$. The failure of the sandwich protocol at $\mu^{2}_{c}$ thus provides a dynamical signature of the quantum critical point, complementary to the static FES diagnostics.
Taken together, these results demonstrate that uMPS–TDVP methods can resolve not only static spectra but also real-time scattering characteristics of an interacting relativistic quantum field theory across a phase transition. The combination of finite-entanglement scaling and real-time dynamics offers a general framework for locating critical points, characterizing quasiparticle content, and assessing the domain of validity of sandwich-geometry scattering constructions. Extensions to larger bond dimensions, refined wave-packet designs, and additional observables (such as detailed phase shifts, inelastic thresholds, and multi-particle production rates) should enable increasingly precise benchmarks of lattice $\phi^{4}$ theory and provide a template for applying tensor-network scattering techniques to more complex quantum field theories.
In future work it will be important to push the simulations to larger bond dimensions and local truncations, and to refine the construction of the incoming and outgoing states beyond the present diagonal-projection scheme, in order to obtain more precise phase shifts and time-delay profiles. The same framework can be extended to compare different scattering bases (e.g. total and relative momentum) and to incorporate improved analytical approximations for the quasiparticle content of the $\phi^{4}$ theory. These developments would turn the present work into the basis of a more comprehensive study of nonperturbative scattering and critical dynamics in interacting quantum field theories.
Acknowledgments
We thank Raghav Jha for sharing his insightful comments on this work in private correspondence, and for his contributions to real-time scattering in Ising Field Theory, which served as a key inspiration for the present study.
W.C. gratefully acknowledges the support provided by Jens Eisert, including financial assistance from Freie Universität Berlin through a research fellowship funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the Research Unit FOR 2724 “Thermische Maschinen in der Quantenwelt" (Fonds 0420611402).
We also thank the Lebanese University for its continuous institutional support.